OpenStudy (raheelafzaal):

y=((2x^1/2)/(2x^1/2+1))^1/2 Find dy/dx plz help me

OpenStudy (anonymous):

What you need to find is the derivative. This is going to be a long and complicated problem. First we can use the power rule and change the problem to 1/2((2x^1/2)/(2x^1/2+1))^-1/2 Does that first step make sense? The power of 1/2 changed to -1/2 to compensate for 1/2 being brought to the front.

OpenStudy (anonymous):

\[y ^{2}=\frac{ 2\sqrt{x} }{ 2\sqrt{x}+1 }\] \[2y \frac{ dy }{ dx }=\frac{ \left( 2\sqrt{x}+1 \right)\frac{ 2 }{ 2\sqrt{x} }-2\sqrt{x}\frac{ 2 }{ 2\sqrt{x} } }{ \left( 2\sqrt{x}+1 \right)^{2}}\] \[ \frac{ dy }{ dx }=\frac{ 1 }{ 2y }*\frac{ 2\sqrt{x}+1-2\sqrt{x} }{ \left( 2\sqrt{x}+1 \right)^{2} }\] \[\frac{ dy }{ dx}=\frac{ 1 }{ 2y \left( 2\sqrt{x}+1 \right)^{2} }\] substitute the value of y and simplify.